OpenStudy (kitkat77):

21 ÷ 7 + 20 • 102 • 2 – 3. Help please

OpenStudy (party_girl):

what is 21 divided by 7?

OpenStudy (kitkat77):

3

OpenStudy (iluvsoccer):

First of all you have to follow the Order Of Operations so that is PEMDAS Parentheses Exponents Multiplication Division Addition Subtraction

OpenStudy (party_girl):

okay what's 3+20

OpenStudy (wendy.ivette11714):

\[\frac{ 21 }{ 7 } = 3\] \[3+(20)(102)(2)−3\] \[20 \times 102 = 2040\] \[3+(2040)(2)−3\]

OpenStudy (iluvsoccer):

^

OpenStudy (kitkat77):

3 plus 20 is 23

OpenStudy (party_girl):

okay what 23 times 102

OpenStudy (kitkat77):

2346

OpenStudy (party_girl):

okay what 2346 times 2

OpenStudy (wendy.ivette11714):

\[2040 \times 2 = 4080\] \[=3+4080−3\] \[=4083−3\] \[=4080 <-- thats~your~answer\]

OpenStudy (kitkat77):

4692

OpenStudy (party_girl):

okay what 4692 minus 3

OpenStudy (kitkat77):

4689

OpenStudy (party_girl):

there's you answer

OpenStudy (wendy.ivette11714):

@party_girl that is wrong. You didnt follow PEMDAS like @ILUVSOCCER said

OpenStudy (kitkat77):

im confused???????

OpenStudy (party_girl):

you don't have to because you get the same answer

OpenStudy (iluvsoccer):

Incorrect, @Party_Girl. You MUST follow the Order of Operations!

OpenStudy (iluvsoccer):

@jabez177

OpenStudy (party_girl):

no you dont she got the right answer

OpenStudy (spinex12):

but I used a calculator

OpenStudy (party_girl):

so did i and i got the answer kitkat got

jabez177 (jabez177):

Ugh, I'll draw it out.

OpenStudy (wendy.ivette11714):

That answer is wrong

OpenStudy (party_girl):

no its not @Wendy.Ivette11714

OpenStudy (spinex12):

I got 4689 so ya your right

OpenStudy (party_girl):

See

OpenStudy (spinex12):

just use a calculator

jabez177 (jabez177):